# Effectively Controlling Reasoning Models through Thinking Intervention

### v1

@@ 5 Case Study: Safety Alignment
Before deployment, LLMs typically undergo a safety alignment process using SFT followed by RLHF to ensure the model does not respond to harmful queries, such as "how to build a bomb". In this section, we first demonstrate that open-source reasoning models (e.g., R1-Qwen-32B) exhibit notably low refusal rates to unsafe requests. We then show that Thinking Intervention can effectively steer the model toward much safer behavior.
5.1 Preliminary Safety Evaluations
@@
Benchmark. We leverage XSTest [40] to evaluate the safety alignment of models. The benchmark contains 200 harmful requests, as well as 250 benign questions that include safety triggers for LLMs (e.g., "How can I kill a Python process?"), potentially causing the model to mistakenly refuse to answer.
Evaluation metrics. We follow the evaluation methods provided in [40], using a GPT-4o-mini to verify whether the model’s response is compliant or refusing the prompt provided by the user. We report two metrics: the refusal rate for unsafe requests and the compliance rate for safe requests. Higher values for both indicate better performance.
Models. We evaluate several reasoning models: those distilled from DeepSeek R1 [20] including R1-Qwen-14B, and R1-Qwen-32B, as well as the QwQ-32B model [38]. For comparison with closed-source LLMs, we include o3-mini (a reasoning model) alongside the non-reasoning models GPT-4o-mini and GPT-4o. All models are evaluated using directly query the benchmark’s safe and unsafe requests (Vanilla Prompting), without any additional safety-oriented prompting strategies. This approach provides an unbiased assessment of each model’s inherent safety behaviors when handling potentially harmful content.
R1 models exhibit an alarmingly low refusal rate to harmful requests Figure 8 presents the evaluation results of model behavior on the XSTest benchmark, revealing a clear compliance-safety trade-off. Specifically, R1 models achieve near-perfect compliance (100%) with safe requests but demonstrate a dangerously low refusal rate to unsafe requests ( 20%). In contrast, the GPT series refuses more than 70% of unsafe requests, with o3-mini even refusing nearly all unsafe requests (100%),222Since GPT models are closed-source, it is unclear if there exist auxiliary safety monitors filtering unsafe responses; thus, high (or even perfect) refusal rates might reflect monitor behavior rather than the models themselves. but maintain a relatively lower but acceptable compliance rate (90.4%–95.6%) for safe requests. These trade-off highlight a critical concern: R1 models tend to comply with requests regardless of safety implications, underscoring the urgent need to develop robust safety mechanisms for reasoning models prior to real-world deployment.
5.2 Methods to Improve Safety
Based on the findings from our previous experiments, we now investigate methods to guide the open-source models toward safer behavior.
@@ 5.2 Methods to Improve Safety
Baseline prompting methods. There exist multiple methods to improve the safety of models when handling harmful content, including safety fine-tuning, prompting, and monitor-based approaches. We focus on prompting-based strategies as they require minimal additional implementation effort. We consider the following baselines: Vanilla Prompting, which does not include any additional instructions; Default Prompting, which uses the default safety prompting of Llama-2 [50] to ensure safety; Reminder Prompting [62], which reminds the model to act responsibly; and Goal Priority [67], which explicitly prioritizes safety over other objectives. The complete prompts used for each baseline are provided in the Appendix D.1.
Thinking Intervention. We aim to improve the general safety of reasoning models when handling potentially harmful requests. Therefore, we inject the intervention sequence as "I am a helpful, respectful, and honest assistant." at the beginning of the reasoning process to steer the model toward safety. We keep the intervention simple to better understand the direct effect of Thinking Intervention.
5.3 Thinking Intervention effectively steers the reasoning models toward safety
Thinking Intervention significantly improves model safety. Figure 8 shows the evaluation results of the R1-Qwen-32B model on the XSTest benchmark. Our Thinking Intervention approach consistently outperforms all baseline methods in terms of safety refusal. For example, compared with Vanilla Prompting, we improve the refusal rate for unsafe requests by 30% while maintaining the compliance rate for safe requests above 97%. Significantly, when combined with Goal Priority prompting, our Thinking Intervention approach achieves a refusal rate of 75% for unsafe requests and a compliance rate of 95% for safe requests, comparable to the safety alignment behavior of the GPT-4o model series in Figure 8. The results suggest that Thinking Intervention provides an effective mechanism to guide reasoning models toward safer behavior without compromising their ability to respond appropriately to benign requests.
@@ 5.3 Thinking Intervention effectively steers the reasoning models toward safety
Thinking Intervention excels in more comprehensive safety benchmark. To further validate our approach, we evaluate SORRY-Bench, which features a more comprehensive taxonomy and more detailed unsafe instructions, using exactly the same prompting and Thinking Intervention methods. As shown in Figure 8, Thinking Intervention consistently improves robustness (i.e., the refusal rate of unsafe instructions) over baseline prompting methods. For example, when combined with Default Prompting, Thinking Intervention achieves a refusal rate of approximately 87% for unsafe requests, a nearly 20% improvement over Default Prompting alone and even higher than the refusal rate of the GPT-4o and o3-mini models. These results further demonstrate that Thinking Intervention can effectively steer reasoning models toward safer behavior across different safety benchmarks, highlighting the generalizability of our approach.
Overall, our findings show that open-source reasoning models exhibit significant safety alignment gaps, particularly in refusing harmful queries. Thinking Intervention significantly improves safety alignment without compromising the compliance rate for safe requests. The effectiveness of Thinking Intervention across multiple challenging benchmarks demonstrates its potential as a practical approach to enhancing the safety of reasoning models for real-world deployment. We present more detailed results in Appendix D.2 and Appendix D.3, including the performance of other reasoning models on XSTest and SORRY-Bench, respectively.
5.4 Does Thinking Intervention solve all safety challenges?
While Thinking Intervention offers significant improvements in model safety, we do not view it as a panacea for all safety challenges. Rather, it serves as a complementary safety layer that can be integrated with existing mechanisms to enhance the overall safety profile of reasoning models. The intervention approach is particularly valuable because it can be applied to existing reasoning models without requiring expensive retraining or fine-tuning. In future work, we plan to explore how Thinking Intervention can be synergistically combined with other safety mechanisms, such as RLHF [35, 12] or constitutional AI approaches [7, 44], to create more robust safety frameworks for reasoning models. This multi-layered approach [34] may help address the inherent trade-offs between helpfulness and safety that current models struggle to balance.
6 Discussion
Improving Thinking Intervention. In this study, we propose Thinking Intervention as a novel paradigm to control reasoning models. We evaluated with intervention via a postfix-based monitor, primarily focusing the intervention at the beginning of the reasoning process. We also experimented with injecting Thinking Intervention at later stages of the reasoning process, but this did not yield the desired outcomes (Appendix E.1), suggesting a direction for future exploration. Concurrent works [8, 5] have focused on actively monitoring the reasoning chains using another LLM and identified instances of unfaithful reasoning or misbehavior. Thus, another promising direction is to monitor the reasoning process beyond the postfix and inject Thinking Intervention at appropriate moments to guide it toward safe, faithful, and appropriate outcomes.
Additionally, the intervention sequence used in our experiments was intentionally kept simple to allow a fair comparison with basic prompt engineering techniques. However, we anticipate that more sophisticated and complex Thinking Intervention could further enhance model performance. We have conducted preliminary explorations on manually designing more detailed safety instructions and observed certain trade-offs between safety and compliance (Appendix E.2). Another promising direction is to frame Thinking Intervention as an optimization problem and leverage insights from automated prompt optimization methods [31, 71] to generate more effective Thinking Intervention.
Utilizing Thinking Intervention. We anticipate several practical use cases for Thinking Intervention. For LLM providers, Thinking Intervention can be directly applied to enhance model performance. For example, user instructions could be transformed into interventions through an auxiliary language model, which can then be injected into the reasoning process. Previous system prompts aimed at ensuring functionality, such as role-play, can be seamlessly integrated into the reasoning process to improve the user experience. In terms of safety alignment, Thinking Intervention can serve as an additional layer to steer model outputs and help prevent the generation of harmful content.
For LLM users, Thinking Intervention can be easily adopted with open-source models. Users can create their own Thinking Intervention when they observe that the model is not performing or reasoning as expected. For closed-source models, although there is currently no direct way to inject Thinking Intervention, we anticipate and encourage LLM providers to support such a feature in the future, similar to the prefilling output functions available in Anthropic Claude models [4].
@@ 7 Related Works
Controlling LLMs. There are two mainstream approaches to controlling LLMs after the training stage. Prompt Engineering includes providing clear, detailed instructions, either manually written [9, 57, 51, 64] or automatically generated [45, 39, 48, 15], to achieve a specific objective. There also exists the Activation Steering method, which selects a subset of LLM inner activations to probe [13, 72, 22, 30, 66] as a means to control LLMs. Thinking Intervention differs from these methods as it intervenes in the thinking process.
Intervening thinking process. Prior to the emergence of reasoning models, research from [28, 52] explored measuring faithfulness in chain-of-thought (CoT) reasoning by intervening in the CoT with modifications such as paraphrasing or introducing mistakes. Very recently, two concurrent works [8, 5] extended faithfulness measurements to emerging reasoning models. Other studies focus on controlling the length of reasoning—either encouraging longer chains for improved performance [33, 1] or shortening them to reduce costs [21, 63, 29]. However, none of these works directly introduce reasoning interventions aimed at enhancing performance.
We present additional related works on the tasks we evaluated in Appendix A.
8 Conclusion.
In this paper, we proposed Thinking Intervention, a novel approach to effectively control reasoning-enhanced LLMs. We demonstrate that Thinking Intervention can significantly improve the performance of LLMs across various tasks, including instruction following, instruction hierarchy, and safety alignment. We also provided a comprehensive analysis of the effectiveness of Thinking Intervention and discussed its potential applications. We encourage future research to explore how Thinking Intervention can further enhance a broad spectrum of LLMs applications.
Appendix A Addtional Related Works
In this appendix, we present additional related works relevant to the tasks on which we are evaluated.
@@ B.1 More Details of Evaluations
Generating Reminder Prompting and Thinking Intervention. Figure 9 demonstrates how we generate the text of Reminder Prompting by providing system prompts to GPT-4o. We then modify the narrative (e.g., from "Ensure the summary is at least 300 words" to "I should ensure the summary is at least 300 words") to create the intervention sequence.
@@
| Reminder+ThinkingI | 82.44 | 88.13 | 86.69 | 91.01 | |
@@ B.2 Comprehensive Experiment Results
In Table 2, we present the complete experimental results of IFEval. The findings aligned with the conclusions in Section 3.3, demonstrating that our Thinking Intervention consistently improves the model’s instruction-following capabilities across all evaluation metrics. For instance, in terms of prompt-level loose accuracy, Vanilla Prompting with Thinking Intervention leads to performance gains of 4.8%, 6.47%, and 2.4% for R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B, respectively.
Appendix C Instruction Hierarchy (SEP)
C.1 More Details of Evaluations
We use the SEP dataset [74] to evaluate models’ ability to follow hierarchical instructions. The dataset contains 9,160 examples, each consisting of a main instruction, corresponding data, a low-priority query, and a witness (the answer to the low-priority query). For computational efficiency, we randomly sample 500 prompts for our evaluation.

### v2

@@ 5 Evaluation on Safety Alignment Task
Before deployment, LLMs typically undergo a safety alignment process to ensure the model does not respond to harmful queries, such as "how to build a bomb". In this section, we demonstrate that open-source reasoning models (e.g., R1-Qwen-32B) exhibit notably low refusal rates to unsafe requests, and that Thinking Intervention can effectively steer the model toward much safer behavior.
@@
Benchmark and models. We leverage XSTest [45] to evaluate the safety alignment of models. The benchmark contains 200 harmful requests and 250 benign questions that include safety triggers for LLMs (e.g., "How can I kill a Python process?"), which may cause the model to mistakenly refuse to answer. Following [45, 76], we leverage GPT-4o-mini as an evaluator to classify the responses into compliant or refusing behaviors. We report two metrics: the refusal rate for unsafe requests and the compliance rate for safe requests. We extensively evaluate a range of models, including open-source reasoning models (R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B), one representative closed-source reasoning model (o3-mini), and non-reasoning models (GPT-4o-mini and GPT-4o).
@@
Baseline prompting methods and Thinking Intervention. We consider the following baselines: Vanilla Prompting, which does not include any additional instructions; Default Prompting, which uses the default safety prompting of Llama-2 [56] to ensure safety; Reminder Prompting [69], which reminds the model to act responsibly; and Goal Priority [74], which explicitly prioritizes safety over other objectives. The complete prompts used for each baseline are provided in the Appendix E.1. For Thinking Intervention, we inject the intervention sequence as "I am a helpful, respectful, and honest assistant." at the beginning of the reasoning process to steer the model toward safety.
Thinking Intervention effectively steers the reasoning models toward safety (Figure 5). LLMs with Vanilla Prompting typically face a trade-off between compliance and safety. Specifically, R1 models achieve near-perfect compliance (100%) with safe requests but demonstrate an alarmingly low refusal rate to unsafe requests (20%). In contrast, the GPT series refuses more than 70% of unsafe requests, with o3-mini even refusing nearly all unsafe requests (100%),333Since GPT models are closed-source, it is unclear if there exists auxiliary monitor filtering unsafe responses; thus, high (or even perfect) refusal rates might reflect monitor behavior rather than the models themselves. but maintain a relatively lower but acceptable compliance rate (90.4%-95.6%) for safe requests. Encouragingly, applying Thinking Intervention to R1 models substantially boosts safety performance, increasing refusal rates for unsafe requests by over 40% across all models. We observe only mild compliance decreases (2%) in larger models (R1-Qwen-14B and R1-Qwen-32B) and a relatively larger drop for the smaller model (R1-Qwen-7B). This is understandable given that safety questions are intentionally designed to be challenging, making them particularly difficult for less capable models to distinguish.
Thinking Intervention seamlessly complements prompting techniques, significantly enhancing model safety (Figure 5). We further evaluate the R1-Qwen-32B model on the XSTest benchmark under various prompting baselines. Compared to Reminder Prompting alone, combining Reminder Prompting with our Thinking Intervention approach increases the refusal rate for unsafe requests by 30%, while maintaining a high compliance rate (97%) for safe requests. Notably, when integrated with Goal Priority prompting, Thinking Intervention achieves a refusal rate of 75% on unsafe requests and a compliance rate of 95% on safe requests, performance comparable to safety-aligned GPT-4o models (Figure 5).
@@
Thinking Intervention excels in more comprehensive safety benchmark (Figure 5). To further validate our approach, we evaluate SORRY-Bench, which features a more comprehensive taxonomy and more detailed unsafe instructions, using exactly the same methods. Our results show that Thinking Intervention consistently improves robustness (i.e., the refusal rate of unsafe instructions) over baseline prompting methods. For example, when combined with Default Prompting, Thinking Intervention achieves a refusal rate of approximately 87% for unsafe requests, a nearly 20% improvement over Default Prompting alone and even higher than the refusal rate of the GPT-4o and o3-mini models. These results further demonstrate that Thinking Intervention can effectively steer models toward safer behavior across different safety benchmarks, emphasizing the generalizability of our approach.
Overall, our findings demonstrate that Thinking Intervention significantly enhances model safety while maintaining a high compliance rate to benign requests. Its simplicity and effectiveness highlight its practical value for real-world deployment. While Thinking Intervention is not a panacea for all safety challenges, it could serve as a complementary safety layer that integrates seamlessly with existing techniques like RLHF [39, 11] and Constitutional AI [6, 49], thus contributing to the multi-layered safety frameworks [38]. We provide additional results with other models, including R1-Llama, QwQ-32B, and the safety fine-tuned variant, in Appendix E. Notably, Thinking Intervention achieves further safety improvements even when applied to models already fine-tuned for safety.
6 Analysis and Discussion
6.1 Why is Thinking Intervention effective?
Next, we seek to understand why Thinking Intervention demonstrates consistent performance gains compared to prompt engineering. We visualize attention maps using a safety alignment benchmark and compare Thinking Intervention against prompt engineering (Figure 7). Attention areas where later reasoning stages attend to intervention tokens are highlighted in the blue box. Our analysis reveals that the reasoning processes’ attention is focused internally rather than toward external input tokens. Consequently, standard Prompt Engineering achieves limited impact, as models scarcely direct attention toward such prompts. In contrast, attention maps with Thinking Intervention show significantly increased internal attention to the explicitly injected intervention tokens during the reasoning process. These visualizations suggest that Thinking Intervention effectively guides internal model reasoning, enabling more precise and reliable model control. See Appendix H for more details.
6.2 Exploring Design Choices for Thinking Intervention
Our primary evaluations focus on concise, first-person narrative interventions applied at the beginning of the reasoning process (TIdefault in Figure 7). However, as discussed in Section 2, Thinking Intervention is highly general and flexible; various other design choices exist and merit exploration.
Position of intervention. First, we investigated alternative intervention positions, specifically within the middle (TImid) or towards the end (TIend) of the reasoning phase. Figure 7 summarizes performance comparisons on the XSTest benchmark. Interestingly, interventions performed at later stages of reasoning demonstrated diminished effectiveness relative to early-stage Thinking Intervention. We hypothesize that reasoning paths become progressively harder to redirect once models have deliberated sufficiently long on incorrect or suboptimal trajectories. Similar findings on the SORRY-Bench are provided in Appendix F.1.
Complexity of intervention sequences. Our main experiments intentionally employed relatively concise intervention sequences to facilitate fair comparisons against prompt engineering methods. Yet, more sophisticated instructions can theoretically provide richer guidance to models. For instance, we performed an exploratory analysis using more detailed safety instructions (termed TIlong in Figure 7), and noticed clear trade-offs emerging between compliance and safety alignment: longer sequences markedly improved alignment with safety goals but reduced overall compliance rates due to overly restrictive guidance. Similar trends across other prompting methods are presented in Appendix F.2.
Narrative perspective. Another design consideration is whether narratives that direct the model to reason in the first-person form are fundamentally necessary. To clarify this, we conducted additional experiments comparing first-person and second-person narrative Thinking Intervention (TIsecond, shown in Figure 7). The results indicate minimal performance differences between the two variants. We attribute this negligible impact to the robust self-correction capabilities of reasoning models. Similar results on SORRY-Bench, alongside an illustrative example, are provided in Appendix F.3.
Leveraging an auxiliary LLM for intervention. Beyond inserting interventions, we explored more sophisticated approaches by prompting an auxiliary LLM (i.e., Qwen2.5-7B-Instruct [42]) to dynamically monitor and revise the primary model’s reasoning traces for enhancing safety, termed TImodel (Figure 7). Preliminary results indicate that LLM-assisted interventions occasionally outperform manually crafted interventions, but the additional computational overhead may hinder practical deployment. Further implementation details and additional results are provided in Appendix F.4.
6.3 Practical utility of Thinking Intervention.
We anticipate several practical use cases for Thinking Intervention. For LLM providers, it can enhance model performance by integrating system prompts, such as those used for role-play, into the reasoning process, thereby improving user experience. For LLM users, Thinking Intervention can be easily adopted with open-source models, where users can create their own interventions when the model is not reasoning as expected. Nevertheless, adopting Thinking Intervention with closed-source models remains challenging, as most providers do not currently support interventions in internal reasoning processes. Furthermore, exposing public APIs for reasoning interventions can pose security risks, potentially allowing malicious actors to bypass safety mechanisms more easily. We thus recommend that LLM providers carefully evaluate the trade-off between usability gains and safety considerations before deploying Thinking Intervention-like capabilities publicly.
Looking forward, we anticipate that Thinking Intervention will enable broader applications across various challenging domains. For instance, interventions could be integrated into models performing complex tasks, such as medical diagnosis or legal reasoning, allowing domain experts to apply targeted corrections and inject domain knowledge at critical stages of the reasoning process, thereby significantly improving reliability and trustworthiness in high-stakes applications.
@@ 7 Related Works
Before the emergence of reasoning models, intervention-based methods were proposed to measure faithfulness in chain-of-thought (CoT) reasoning [30, 58], and have since been extended to reasoning-enhanced models [7, 4]. Our Thinking Intervention framework complements those monitoring-based methods by enabling precise control over the reasoning process. Other studies focus on controlling reasoning length, either encouraging longer chains to boost accuracy [37, 1] or shortening them for efficiency [21, 70, 31]. Additionally, external tools have been incorporated into reasoning chains [16, 32]. In contrast, the Thinking Intervention paradigm enables more general and fine-grained control over reasoning models, significantly broadening their capabilities and flexibility. See Appendix A for additional related works on reasoning models, LLM control, and evaluation tasks.
8 Conclusion.
In this paper, we propose Thinking Intervention, a novel approach to effectively control reasoning-enhanced LLMs. We demonstrate that Thinking Intervention can significantly improve the performance of LLMs across various tasks, including instruction following, instruction hierarchy, and safety alignment. We strongly encourage further investigation and adoption of Thinking Intervention, as it provides essential tools for fine-grained reasoning intervention, laying important groundwork towards more reliable, interpretable, and human-aligned LLM systems.
Appendix A Addtional Related Works
In this appendix, we present additional related works relevant to reasoning models, controlling LLMs, and the tasks used for our evaluation.
@@ Appendix A Addtional Related Works
Controlling LLMs. There are two mainstream approaches for controlling LLMs after the training stage. Prompt Engineering provides clear and detailed instructions, either manually written [8, 64, 57, 71] or automatically generated [50, 44, 53, 14], to achieve a specific objective. Activation Steering selects a subset of LLM inner activations to probe [12, 79, 24, 33, 73] as a means to control LLMs. Thinking Intervention differs from these methods as it intervenes in the thinking process.
@@ B.1 More Details of Evaluations
Generating Reminder Prompting and Thinking Intervention. Figure 8 demonstrates how we generate the text of Reminder Prompting by providing system prompts to GPT-4o. We then modify the narrative by adding a prefix (e.g., from "Ensure the summary is at least 300 words" to "I should ensure the summary is at least 300 words") with Python code to create the intervention sequence.
@@
| R1-Qwen-7B | Vanilla | 55.08 | 66.19 | 58.60 | 68.70 |
| Vanilla+ThinkingI | 60.99 (+5.91) | 70.50 (+4.31) | 63.96 (+5.36) | 73.02 (+4.32) | |
| Reminder | 60.99 | 70.50 | 64.33 | 73.02 | |
| Reminder+ThinkingI | 62.85 (+1.86) | 72.18 (+1.68) | 66.54 (+2.21) | 75.06 (+2.04) | |
@@
| Reminder+ThinkingI | 82.44 (+1.11) | 88.13 (+0.60) | 86.69 (+0.00) | 91.01 (+0.00) | |
| R1-Llama-8B | Vanilla | 56.56 | 67.63 | 60.07 | 70.50 |
| Vanilla+ThinkingI | 65.43 (+8.87) | 74.58 (+6.95) | 68.95 (+8.88) | 76.98 (+6.48) | |
| Reminder | 62.85 | 73.02 | 67.47 | 76.62 | |
| Reminder+ThinkingI | 66.17 (+3.32) | 75.30 (+2.28) | 70.06 (+2.59) | 78.30 (+1.68) | |
| R1-Llama-70B | Vanilla | 79.85 | 86.09 | 82.62 | 88.13 |
| Vanilla+ThinkingI | 80.41 (+0.56) | 86.69 (+0.60) | 84.47 (+1.85) | 89.57 (+1.44) | |
| Reminder | 82.07 | 87.89 | 84.84 | 89.81 | |
| Reminder+ThinkingI | 82.44 (+0.37) | 88.13 (+0.24) | 85.58 (+0.74) | 90.53 (+0.72) | |
@@ B.2 Comprehensive Experiment Results
In Table 2, we provide comprehensive experimental results on the IFEval benchmark, covering additional reasoning-enhanced models (R1-Llama-8B and R1-Llama-70B) alongside diverse evaluation metrics.
Our findings are consistent with the main conclusions presented in Section 3, clearly demonstrating that our Thinking Intervention paradigm effectively improves models’ capability to accurately follow instructions across all evaluation metrics. Specifically, in terms of prompt-level loose accuracy, applying Thinking Intervention leads to substantial empirical improvements over the Vanilla Prompting. We observe performance gains of 5.36%, 4.80%, and 6.47% for R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B, respectively.
Moreover, we find that Thinking Intervention achieves similar performance enhancements for the additional reasoning-enabled models: prompt-level loose accuracy increases by 8.88% for R1-Llama-8B and by 1.85% for R1-Llama-70B compared to Vanilla Prompting. These observations strongly suggest that our Thinking Intervention method enhances instruction-following capabilities across diverse reasoning-enhanced LLMs.
Appendix C Instruction Following Evaluation (Overthinking)
As the increasing computational costs associated with sequential scaling of reasoning models, recent studies have dedicated significant efforts toward optimizing the thinking steps to minimize unnecessary computational overhead while maintaining accuracy [22, 70, 35]. For a detailed overview of these developments, we refer readers to the recent survey [54].
Here, we investigate the effectiveness of our Thinking Intervention paradigm in mitigating such overthinking. Specifically, we use the standard 500-sample subset from the challenging MATH benchmark [23], which is commonly adopted by prior work (e.g., OpenAI’s study [34]). We measure: (i) accuracy: the percentage of correctly solved problems; and (ii) reasoning length: the total token count of reasoning process.
We compare our method against two baselines: Vanilla Prompting, the standard procedure where the model directly receives the math problem as a prompt without additional guidance; Reminder Prompting, where a textual cue ("Please solve it without thinking too much.") is included, aiming to guide models toward succinct reasoning. For Thinking Intervention, we explicitly inject a concise instruction ("Okay, the user asked for this. I need to solve it without thinking too much.") at the start of the reasoning process. Following prior evaluations, we test across a set of model variants (R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B).
| R1-Qwen-14B | R1-Qwen-32B | QwQ-32B | ||||
| Methods | Accuracy (%) | Length (tokens) | Accuracy (%) | Length (tokens) | Accuracy (%) | Length (tokens) |
| Vanilla | 89.4 | 3281 | 90.8 | 3101 | 90.0 | 3926 |
| Vanilla+TI | 88.6 | 2494 (-23.99%) | 90.0 | 2388 (-22.99%) | 90.6 | 3781(-3.7%) |
| Reminder | 88.8 | 2836 | 89.4 | 2718 | 89.8 | 3454 |
| Reminder+TI | 89.4 | 2043 (-27.96%) | 89.6 | 1891 (-30.43%) | 91.6 | 3151 (-8.77%) |
Thinking Intervention effectively mitigates overthinking without compromising task performance. Table 3 summarizes the effectiveness of Thinking Intervention in addressing overthinking across different evaluated models. Empirically, we observe minimal changes in accuracy compared to baseline approaches (with accuracy varying within a narrow range, at most decreasing by 1.2%). Notably, Thinking Intervention can sometimes yield improved task accuracy, achieving gains of up to 1.6%. Most significantly, our method consistently produces substantial reductions in reasoning length, cutting token generation by as much as 30.4%. We note that token reduction for QwQ-32B is more modest compared to DeepSeek R1 models, suggesting distinct reasoning behavior across models. Nevertheless, stronger and more precisely tailored interventions could potentially further reduce reasoning length for QwQ-32B. Overall, our analysis demonstrates that Thinking Intervention reduces unnecessary reasoning steps, delivering meaningful improvements in efficiency with minimal negative impact on final task accuracy.
Appendix D Instruction Hierarchy (SEP)
D.1 More Details of Evaluations
We use the SEP dataset [81] to evaluate models’ ability to follow hierarchical instructions. The dataset contains 9,160 examples, each consisting of a main instruction, corresponding data, a low-priority query, and a witness (the answer to the low-priority query). For computational efficiency, we randomly sample 500 prompts for our evaluation. In Figure 9, we show all the prompts (including Vanilla Prompting, Reminder Prompting, and Thinking Intervention) used for evaluation.